40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4806     Entry time: Fri Jun 10 18:49:40 2011
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: Intensity Noise after the MC 

Last night the relative intensity noise (RIN) of the beam after MC was measured.

It looks like the RIN is dominated by the motion of the MC mirrors, possibly the angular motions because we don't have any angular stabilization servos.

Suresh will estimate the contribution from the MC mirrors' angular motion to the RIN in order to see if this plot makes sense.

 

(RIN)

 The spectrum below 30 Hz seems to be dominated by the motion of the MC suspensions because it very resembles the spectra of those.

Above 30 Hz the spectrum becomes somehow flat, which I don't know why at the moment.

A rough estimation of the shot noise gave me a level of 10^{-9} RIN/sqrtHz, which is way below the measured spectrum.

RIN_afterMC.png

 

(Setup)

 All of the suspended mirrors were intentionally misaligned except for the MC mirrors and PRM to avoid interference from the other optics.

In this setup it allows us to measure the intensity noise of the laser which is transmitted from MC.

The beam transmitted from MC is reflected by PRM and finally enters into the REFL11 RFPD.

The DC signal from the RFPD was acquired at C1:LSC-REFL_DC_DQ as the laser intensity.

As well as the RIN measurement I took a spectrum when the beam is blocked by a mechanical shutter on the PSL table.

This data contains the dark noise from REFL11 and circuit noise from the whitening, AA board and ADC. It is drawn in black in the plot.

The cut off at 15 Hz is possibly due to the digital unwhitening (two poles at 15 Hz and two zeros at 150 Hz) filter to correct the analog whitening filter.

ELOG V3.1.3-